Owen’s Electronics has nine operating plants in seven southwestern states. Sales for last year were $100 million, and the balance sheet at year-end is similar in percentage of sales to that of previous years (and this will continue in the future). All assets (including fixed assets) and current liabilities will vary directly with sales. The firm is working at full capacity. Balance Sheet (in $ millions) Assets Liabilities and Stockholders' Equity Cash $ 7 Accounts payable $ 20 Accounts receivable 25 Accrued wages 7 Inventory 28 Accrued taxes 13 Current assets $ 60 Current liabilities $ 40 Fixed assets 45 Notes payable 15 Common stock 20 Retained earnings 30 Total assets $ 105 Total liabilities and stockholders' equity $ 105 Owen’s Electronics has an aftertax profit margin of 10 percent and a dividend payout ratio of 45 percent. If sales grow by 20 percent next year, determine how many dollars of new funds are needed to finance the growth. (Do not round intermediate calculations. Enter your answer in dollars, not millions, (e.g., $1,234,567).)
Answer #1

Answer:

Sales last year = $100 million = $100,000,000

Projected Sales next year = $100,000,000 * (1 + 20%) = $120,000,000

∆ Sales = 120000000 - 100000000 = $20,000,000

Assets = $105 million = $105,000,000

Spontaneous liabilities = Current liabilities = $40 million = $40,000,000

AFN = (Assets / Sales) * ∆ Sales - (Spontaneous liabilities / Sales) * ∆ Sales - Projected Sales * After-tax profit margin * (1 - Dividend Payout ratio)

= (105000000 / 100000000) * 20000000 - (40000000 / 100000000) * 20000000 - 120000000 * 10% * (1 - 45%)

= $6,400,000

New funds needed = $6,400,000